# On the number of modes of Gaussian kernel density estimators

### v2

@@ 1.1 Setup and main result
In this paper we are interested in determining the expected number of modes (local maxima) of $\widehat{P}_{n}$ over $\mathbb{R}$. While this is a classical question, addressed in even more general settings than (1.1)—such as non-Gaussian kernels, compactly supported samples, and higher dimensions [MMF92, Mam95, KM97]—a definite answer has not been given in the literature. Indeed, the best-known results fall into one of two settings: either considering samples drawn from a compactly supported density (instead of $N(0,1)$ as done here), or counting the modes within a fixed compact interval. In the special case of the Gaussian KDE (1.1), one has in the latter setting for instance
Theorem 1.1 ([Mam95, Thm. 1]).
Let $\widehat{P}_{n}$ be the Gaussian KDE defined in (1.1), with bandwidth $h>0$, of $X_{1},\dots,X_{n}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}N(0,1)$. Asymptotically as $n\to\infty$, the expected number of modes of $\widehat{P}_{n}$ in a fixed interval $[a,b]$ is $1+o(1)$ if $1\ll\beta\ll n^{\frac{2}{3}}$, and $\widetilde{\Theta}\mathopen{}\mathclose{{}\left(\sqrt{\beta}}\right)$ if $n^{\frac{2}{3}}\lesssim\beta\ll n^{2}/\log^{6}n$.
In [MMF92, Mam95, KM97], the authors additionally conduct more refined casework on the bandwidth to provide more precise estimates, such as pinpointing the leading constants. In fact, [MMF92] does count modes in $\mathbb{R}$, but the underlying distribution of the samples $X_{i}$ is supported on a closed interval (thus excluding $N(0,1)$), so there are no modes outside the interval anyway.
Our main result provides the answer to the case of counting modes of (1.1) over $\mathbb{R}$, and reads as follows. In particular, it generalizes Theorem 1.1.
@@
Figure 1: A realization of the random function (1.1) for $n=10^{4}$, with $\beta=100$ (left) and $\beta=300$ (right).
Several comments are in order.
@@
  • To better appreciate the range of values for $\beta$ in this theorem as well as subsequent ones, we use minimax theory as a benchmark; see, e.g., [Tsy09]. The reparametrization $h=\beta^{-\frac{1}{2}}$ is motivated by the connection to the Transformer model described in Section 1.2. Using an optimal bias-variance tradeoff [Tsy09, Chapter 1], we see that the optimal scaling of the bandwidth parameter $h$ depends on the smoothness of the underlying density of interest: if the underlying density has $s$ bounded (fractional) derivatives, then the optimal choice of $h$ is given by $h\asymp n^{-\frac{1}{2s+1}}$. This gives $\beta\asymp n^{\frac{2}{2s+1}}$. For $s>0$, we get $\beta\in[n^{c},n^{2-c}]$ for some $c>0$. In particular, the transition of the number of modes from 1 to $\sqrt{\beta}$ in Theorem 1.1 is achieved for $\beta\approx n^{\frac{2}{3}}$, which is the optimal choice for Lipschitz densities. The message of our main Theorem 1.2 below is that this scaling in $\sqrt{\beta}$ is the prevailing one for the whole range $\beta\in[n^{c},n^{2-c}]$ if one does not restrict counting modes in a bounded interval $[a,b]$.
  • Point 2. in Theorem 1.2 shows that most of the modes are at distance at least $C\log n$ from the origin provided $\beta>n^{\frac{2-C}{3}}$ for $C>0$ small. This corresponds to a choice of a bandwidth adapted to smoothness $s<1$. This result is in agreement with and completes the picture drawn by Theorem 1.1.
@@
Through refined computations, one can determine the modes in the regime $1\ll\beta\ll n^{2}$ and also pinpoint the leading constant. For the sake of clarity, we stick to the regime where
$2\log n-\log\beta\asymp\log\beta\asymp\log n,$
and comment on how to do expand the regime in appropriate places.
@@
We further motivate Point 2. in Theorem 1.2 by considering a qualitative picture of the distribution of the modes displayed in Figure 4.
  • Near the origin, we find most of the samples $X_{i}$ and they are densely packed in the shape of a Gaussian. The corresponding Gaussian summands in 1.1 cancel to create one mode, as shown already in Theorem 1.1.
  • In the two intervals of length $\Theta\mathopen{}\mathclose{{}\left(\sqrt{\log\beta}}\right)$, the samples $X_{i}$ are separated enough that the corresponding Gaussian summands do not cancel, but rather form $\Omega\mathopen{}\mathclose{{}\left(\sqrt{\log\beta}}\right)$ isolated bumps, as discussed in more generality in [DG85, Section 9.3].
We revisit this discussion and Figure 4 in Remark 3.3.
To prove Theorem 1.2, we truncate the real line to the interval
@@
and so $T$ is well-defined for large enough $\beta$. Motivated by Theorem 1.2, we also define the interval
@@
if $\beta\leqslant n^{\frac{2}{3}}$ and define $T^{\prime}=\varnothing$ if $\beta>n^{\frac{2}{3}}$.
We see how Theorem 1.2 implies Theorem 1.1 with $h=\beta^{-\frac{1}{2}}$. From the former, we see that almost all the modes lie in $T\setminus T^{\prime}$. If $h\gg n^{-\frac{1}{3}}$ so that $\sqrt{2\log n-3\log\beta}\gg 1$, then $T\setminus T^{\prime}$ is disjoint from $[a,b]$, so there are few modes in $[a,b]$; if $h\ll n^{-\frac{1}{3}}$ so $\sqrt{2\log n-3\log\beta}\ll 1$, the fixed interval $[a,b]$ contains a near constant fraction of length of $T\setminus T^{\prime}$ and thus a near constant fraction of the modes.
Figure 2: (Left) Plot of the average number of modes as a function of $\beta$ for $n=10^{3}$ (top) and $n=10^{4}$ (bottom). (Right) Log-log plot for $n=10^{3}$ (top) and $n=10^{4}$ (bottom); the predicted linear regression line (red) corroborates a power-law of the form average # of $\text{average \# of modes}\approx 0.179\cdot\beta^{0.504}$, in line with Theorem 1.2.
1.2 Motivation
The question of estimating the number of modes as a function of the bandwidth has a plethora of applications in statistical inference and multimodality tests—see [MMF92, Mam95, KM97] and the references therein. Another application which has stimulated some of the recent progress on the topic is data clustering. The latter can be achieved nonparametrically using a KDE, whose modes, and hence clusters, can be detected using the mean-shift algorithm [FH75, Che95, CM02, CP00, CPW03, CP07, RL14, CP15], which can essentially be seen as iterative local averaging. The main idea in mean-shift clustering is to perform a mean-shift iteration starting from each data point and then define each mode as a cluster, with all points converging to the same mode grouped into the same cluster. The analysis of this algorithm has led to upper bounds on the number of modes of (1.1) [CPW03].
We were instead brought to this problem from another perspective, motivated by the study of self-attention dynamics [SABP22, GLPR23, GLPR24, GRRB24]—a toy model for Transformers, the deep neural network architecture that has driven the success of large language models [VSP+17]. These dynamics form a mean-field interacting particle system
It is shown that for almost every initial configuration $x_{1}(0),\ldots,x_{n}(0)$, and for $\beta\geqslant 0$ in dimension $\geqslant 3$, or $\beta\leqslant 1\vee\beta\gtrsim n^{2}$ in dimension $2$, all particles converge to a single cluster in infinite time [GLPR23]. Rather than converging quickly, the authors in [GKPR24] prove that the dynamics instead manifest metastability: particles quickly approach a few clusters, remain in the vicinity of these clusters for a very long period, and eventually coalesce into a single cluster in infinite time. Concurrently, and using different methods, the authors in [BPA24] show a similar result: starting from a perturbation of the uniform distribution, beyond a certain time, the empirical measure of the $n$ particles approaches an empirical measure of $O(\sqrt{\beta})$-equidistributed points on the circle in the mean-field limit, and stays near it for long time. This is done by a study of the linearized system and leveraging nonlinear stability results from [Gre00].
Our interest lies in counting the number of clusters during the first metastable phase in dimension $d=2$. At time $\tau=0$, particles are initialized at $n$ iid points from the uniform distribution on the circle. In turn, the stationary points of $\mathsf{K}_{\beta^{-1/2}}\ast\mu_{0}$ partition the circle into intervals, with points clustering within their respective interval. This highlights the importance of counting the number of stationary points.
Here, we focus on a simplified setting by working on the real line instead of the circle (or higher-dimensional spheres), but we believe the analysis could be extended to these cases pending technical adaptations. Notwithstanding, Theorem 1.2 reflects what is seen in simulations (Figure 3).
1.3 Sketch of the proof
The spirit of the proof of results such as Theorem 1.1 and others presented in [MMF92, Mam95, KM97] is similar to ours—one applies the Kac-Rice formula (Theorem 2.1) to a Gaussian approximation of $(\widehat{P}_{n}^{\prime}(t),\widehat{P}_{n}^{{}^{\prime\prime}}(t))$ and argues its validity. However, the main limitation of these works is that modes are counted in a fixed and finite interval $[a,b]$ (and $[0,1]^{d}$ in the higher dimensional cases). Extending these techniques to the whole real line demands for different, significantly stronger, approximation results using Edgeworth expansions.
We sketch the key ideas that allow us to count modes over $\mathbb{R}$. We use the Kac-Rice formula to compute the expected number of modes of $\widehat{P}_{n}$ in the symmetric interval $T$ and $T^{\prime}$ defined in 1.2 and 1.3. All asymptotics are as $n,\beta\to\infty$.
@@ 1.3 Sketch of the proof
    In expectation over $X_{i}$, the number of modes of $\widehat{P}_{n}$ in $T^{\prime}$ is $O\mathopen{}\mathclose{{}\left(e^{-\frac{\omega(\beta)}{4}}\sqrt{\beta\log\beta}}\right)$.
The Kac-Rice computation appears tractable only when the joint distribution of $(\widehat{P}_{n}^{\prime}(t),\widehat{P}_{n}^{\prime\prime}(t))$ is Gaussian, which it is not. To overcome this obstacle, we apply the Kac-Rice formula over a Gaussian approximation of the joint distribution in Section 2. For the specific underlying density and KDE in 1.1, we are able to justify in Section 3 the approximation for all $t$ in the growing interval $T$ instead of a fixed interval. This is why Theorem 1.1 only counts modes in a fixed interval.
To show the validity of the Gaussian approximation, we use the Edgeworth expansion of the joint distribution of $(\widehat{P}_{n}^{\prime}(t),\widehat{P}_{n}^{\prime\prime}(t))$ around the Gaussian distribution with matching first two moments. We bound the error due to the third order term of the expansion directly, and deal with the higher order terms by appealing to the error bounds of densities in the Edgeworth approximation similar to [BR10, Theorem 19.2]. This strategy has been used in [BCP19], in a different context. We note that [KM97] employ the same theorem to justify the Gaussian process approximation over a fixed interval.
In doing so, we will see that the normal approximation is invalid outside of $T$ (see Remark 3.3), but crucially $T$ is sufficiently large to cover almost all modes, as observed empirically in Remark 1.5 and Figure 4 and given below.
@@
If $n^{c}\lesssim\beta\lesssim n^{2-c}$ for arbitrarily small $c>0$, then the expectation over $X_{i}$ of the number of modes of $\widehat{P}_{n}$ that lie outside of $T$ is $O\mathopen{}\mathclose{{}\left(\sqrt{\beta\exp(\omega(\beta))}}\right)$.
We prove this in Section 4.2 with an argument from scale-space theory: we bound the number of modes outside $T$ by the number of samples $X_{i}$ outside $T$, which we then bound naively. This is precisely the argument used by [CPW03, Theorem 2] to show Gaussian mixtures over $\mathbb{R}$ with $n$ components must have at most $n$ modes. This argument crucially relies on the kernel density estimator being Gaussian (see Remark 4.4). Now, Theorem 1.2 follows from Propositions 1.6 and 1.7 upon recalling that $1\ll\omega(\beta)\ll\log\log\beta$.

### v3

@@ 1.1 Setup and main result
In this paper we are interested in determining the expected number of modes (local maxima) of $\widehat{P}_{n}$ over $\mathbb{R}$. While this is a classical question, addressed in even more general settings than (1.1)—such as non-Gaussian kernels, compactly supported samples, and higher dimensions [mammen91, mammen95, mammen97]—a definite answer has not been given in the literature. Indeed, the best-known results fall into one of two settings: either considering samples drawn from a compactly supported density (instead of $N(0,1)$ as done here), or counting the modes within a fixed compact interval. In the special case of the Gaussian KDE (1.1), one has in the latter setting for instance
Theorem 1.1 ([mammen95, Thm. 1]).
Let $\widehat{P}_{n}$ be the Gaussian KDE defined in (1.1), with bandwidth $h\coloneqq\beta^{-\frac{1}{2}}>0$, of $X_{1},\dots,X_{n}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}N(0,1)$. Asymptotically as $n\to\infty$, the expected number $N$ of modes of $\widehat{P}_{n}$ in a fixed interval $[a,b]$ is
  • $\mathbf{1}\{0\in[a,b]\}+o(1)$ if $\beta\ll n^{\frac{2}{5}}$,
  • $\Theta(1)$ if $\beta\asymp n^{\frac{2}{5}}$
  • $\Theta\mathopen{}\mathclose{{\left(n^{-\frac{1}{2}}\beta^{\frac{5}{4}}}}\right)=o\mathopen{}\mathclose{{\left(\sqrt{\beta}}}\right)$ if $n^{\frac{2}{5}}\ll\beta\ll n^{\frac{2}{3}}$,
  • and $\Theta\mathopen{}\mathclose{{\left(\sqrt{\beta}}}\right)$ if $n^{\frac{2}{3}}\lesssim\beta\ll n^{2}/\log^{6}n$.
In [mammen91, mammen95, mammen97], the authors additionally conduct more refined casework on the bandwidth to provide more precise estimates, such as pinpointing the leading constants. In fact, [mammen91] does count modes in $\mathbb{R}$, but the underlying distribution of the samples $X_{i}$ is supported on a closed interval (thus excluding $N(0,1)$), so there are no modes outside the interval anyway.
In the case of counting modes of (1.1) over $\mathbb{R}$, [carreira2003number] provides an upper bound of $n$ using scale-space theory by showing adding components one-by-one to a Gaussian mixture increases the mode-count by at most one each time. Our main result velow provides a precise answer in this case. For the sake of clarity, we stick to the regime where
$2\log n-\log\beta\asymp\log\beta\asymp\log n.$
Through refined computations, one can determine the modes in the regime $1\ll\beta\lesssim n^{2}/\log^{\Theta(1)}(n)$ and also pinpoint the leading constant. We state our main theorem, and will comment on how to do expand the regime in appropriate places.
@@
Figure 1: A realization of the kernel density estimator $\widehat{P}_{n}$ in (1.1) for $n=10^{4}$, with $\beta=100$ (left) and $\beta=300$ (right). Larger $\beta$ narrows the Gaussian kernel, which sharpens $\widehat{P}_{n}$ and reveals more small peaks on the shoulders, while the central peak remains single. Theorem˜1.2 later quantifies where and how many such peaks appear.
In fact, we can bound on the rate of convergence of the little-$o$ in Point 2. This is spelled-out in Propositions˜1.7 and 1.8. Several comments are in order.
@@
  • To better appreciate the range of values for $\beta$ in this theorem as well as subsequent ones, we use minimax theory as a benchmark; see, e.g., [Tsy09]. The reparametrization $h=\beta^{-\frac{1}{2}}$ is motivated by the connection to the Transformer model described in Section˜1.2. Using an optimal bias-variance tradeoff [Tsy09, Chapter 1] [^1] With $h=\beta^{-\frac{1}{2}}$, the usual bias–variance calculus for $s$-smooth densities gives $h\asymp n^{-\frac{1}{2s+1}}$ and hence $\beta\asymp n^{\frac{2}{2s+1}}$  [Tsy09, Ch. 1]., we see that the optimal scaling of the bandwidth parameter $h$ depends on the smoothness of the underlying density of interest: if the underlying density has $s$ bounded (fractional) derivatives, then the optimal choice of $h$ is given by $h\asymp n^{-\frac{1}{2s+1}}$. This gives $\beta\asymp n^{\frac{2}{2s+1}}$. For $s>0$, we get $\beta\in[n^{c},n^{2-c}]$ for some $c>0$. In particular, the transition of the number of modes from 1 to $\sqrt{\beta}$ in Theorem˜1.1 is achieved for $\beta\approx n^{\frac{2}{3}}$, which is the optimal choice for Lipschitz densities. The message of our main Theorem 1.2 above is that this scaling in $\sqrt{\beta}$ is the prevailing one for the whole range $\beta\in[n^{c},n^{2-c}]$ if one does not restrict counting modes in a bounded interval $[a,b]$.
  • Point 2. in Theorem˜1.2 shows that most of the modes are at distance at least $C\log n$ from the origin provided $\beta>n^{\frac{2-C}{3}}$ for $C>0$ small. This corresponds to a choice of a bandwidth adapted to smoothness $s<1$. This result is in agreement with and completes the picture drawn by Theorem˜1.1.
@@
We further motivate Point 2. in Theorem˜1.2 by considering a qualitative picture of the distribution of the modes displayed in Figure˜4.
  • Near the origin, we find most of the samples $X_{i}$ and they are densely packed in the shape of a Gaussian. The corresponding Gaussian summands in ˜1.1 cancel to create one mode, as shown already in Theorem˜1.1.
  • In the two intervals of length $\Theta\mathopen{}\mathclose{{\left(\sqrt{\log\beta}}}\right)$, the samples $X_{i}$ are separated enough that the corresponding Gaussian summands do not cancel, but rather form $\Omega\mathopen{}\mathclose{{\left(\sqrt{\log\beta}}}\right)$ isolated bumps, as discussed in more generality in [devgyo, Section 9.3].
Write $\widehat{P}_{n}^{\prime}(t)=\mathbb{E}\,\widehat{P}_{n}^{\prime}(t)+(\widehat{P}_{n}^{\prime}(t)-\mathbb{E}\widehat{P}_{n}^{\prime}(t))$. The first term (“bias”) reflects the deterministic drift toward a single broad mode, while the second (“variance”) creates random sign changes that generate extra modes. For the rescaled field $F_{n}(t)=-c\,\widehat{P}_{n}^{\prime}(t)$ in (2.2), Lemma˜2.2 yields
$\mathrm{SNR}(t)^{2}\coloneqq\frac{|\mathbb{E}F_{n}(t)|^{2}}{\mathrm{Var}F_{n}(t)}\asymp nt^{2}\beta^{-\frac{3}{2}}e^{-\frac{t^{2}}{2}}.$
When $\mathrm{SNR}(t)\gg 1$ the bias dominates and no additional modes appear; when $\mathrm{SNR}(t)\ll 1$ the variance takes over and modes proliferate. The crossover $\mathrm{SNR}(t)\approx 1$ gives the inner edge $t^{2}\approx 2\log n-3\log\beta$ (up to $\log t$ terms). On the other hand, to form isolated bumps we also need at least one point in a kernel window of width $h=\beta^{-\frac{1}{2}}$, i.e. $n\varphi(t)h\approx 1$, which gives the outer edge $t^{2}\approx 2\log n-\log\beta$. Together these two thresholds select the belt $t^{2}\in[2\log n-3\log\beta,2\log n-\log\beta]$ and, in particular, center the localization at $|t|\approx\sqrt{2\log n-\log\beta}$.
We revisit this discussion and Figure˜4 in Remark˜3.4.
Remark 1.5 (Belt width).
From Section˜2.3 the Kac–Rice density is proportional to $\sqrt{\beta}e^{-A_{t}}$ with $A_{t}\asymp\beta^{-\frac{3}{2}}nt^{2}e^{-\frac{t^{2}}{2}}$, so the mass concentrates where $A_{t}=O(1)$. This pins down $t^{2}=2\log n-c\log\beta+O(\log\log\beta)$ with $c\in\{1,3\}$—precisely the endpoints of Theorem˜1.2—and converting from $t^{2}$ to $t$ turns the $2\log\beta$ gap into a belt of length $\Delta t\approx(2t_{*})^{-1}\cdot 2\log\beta\asymp\sqrt{\log\beta}$ around $t_{*}\approx\sqrt{2\log n-\log\beta}$.
@@
We compare our result with Theorem˜1.1. Let $J$ be the union of the symmetric intervals of length $\Theta(\sqrt{\log\beta})$ in Point 2, i.e. the “two belts” region where almost all modes lie. Our proof will show that the density of modes is $\Theta(\sqrt{\beta})$ whenever $t\in J$, and $o(\sqrt{\beta})$ whenever $t\not\in J$. This gives the Theorem˜1.2 upon integrating over $t$, and explains the threshold $\beta\asymp n^{\frac{2}{3}}$ in Theorem˜1.1:
  • If $\beta\gtrsim n^{\frac{2}{3}}$, then $a,b]\subset J$ for sufficiently large $n$ and $\beta$, so the density of modes is $\Theta(\sqrt{\beta})$ everywhere on $[a,b]$, giving $\Theta(\sqrt{\beta})$ modes in total.
  • If $\beta\ll n^{\frac{2}{3}}$, then $[a,b]$ is between (and outside of) the two belts of $J$ for sufficiently large $n$ and $\beta$, so the density of modes is $o(\sqrt{\beta})$ everywhere on $[a,b]$, giving $o(\sqrt{\beta})$ modes in total. In fact, Proposition˜1.7 shows this symmetric interval $T^{\prime}$ between the two belts has $O(\sqrt{\beta})$ modes in total. We can see that $T^{\prime}$ has length $\omega_{n\to\infty}(1)$ and the mode density is increasing as we move away from 0, so the number of modes in $[a,b]$ must be a $o(1)$-fraction of the modes in $T^{\prime}$, i.e. it is $o(\sqrt{\beta})$.
Hence, this corollary of our result implies the last two bullet points of Theorem˜1.1. Similarly, by truncating to more refined intervals separated by $t^{2}=2\log n-5\log\beta$, we can hope to recover the threshold $\beta\asymp n^{\frac{2}{5}}$ given in the first three bullet points of Theorem˜1.1, but we do not pursue this here.
Figure 2: (Left) Plot of the average number of modes as a function of $\beta$ for $n=10^{3}$ (top) and $n=10^{4}$ (bottom). (Right) Log-log plot for $n=10^{3}$ (top) and $n=10^{4}$ (bottom); the predicted linear regression line (red) corroborates a power-law of the form $\text{average \# of modes}\approx 0.179\cdot\beta^{0.504}$, in line with Theorem˜1.2.
1.2 Motivation
The question of estimating the number of modes as a function of the bandwidth has a plethora of applications in statistical inference and multimodality tests—see [mammen91, mammen95, mammen97] and the references therein. Another application which has stimulated some of the recent progress on the topic is data clustering. The latter can be achieved nonparametrically using a KDE, whose modes, and hence clusters, can be detected using the mean-shift algorithm [fukunaga1975estimation, cheng1995mean, comaniciu2002mean, carreira2000mode, carreira2003number, carreira2007gaussian, rodriguez2014clustering, carreira2015review], which can essentially be seen as iterative local averaging. The main idea in mean-shift clustering is to perform a mean-shift iteration starting from each data point and then define each mode as a cluster, with all points converging to the same mode grouped into the same cluster. The analysis of this algorithm has led to upper bounds on the number of modes of (1.1) [carreira2003number].
We were instead brought to this problem from another perspective, motivated by the study of self-attention dynamics [sander2022sinkformers, geshkovski2023mathematical, geshkovski2024emergence, geshkovski2024measure]—a toy model for Transformers, the deep neural network architecture that has driven the success of large language models [vaswani2017attention]. These dynamics form a mean-field interacting particle system
It is shown that for almost every initial configuration $x_{1}(0),\ldots,x_{n}(0)$, and for $\beta\geqslant 0$, all particles converge to a single cluster in infinite time [geshkovski2023mathematical, criscitiello2024synchronization, polyanskiy2025synchronization]. Rather than converging quickly, [geshkovski2024dynamic] prove that the dynamics instead manifest metastability: particles quickly approach a few clusters, remain in the vicinity of these clusters for a very long period, and eventually coalesce into a single cluster in infinite time. Concurrently, and using different methods, [bruno2024emergence] show a similar result: starting from a perturbation of the uniform distribution, beyond a certain time, the empirical measure of the $n$ particles approaches an empirical measure of $O(\sqrt{\beta})$-equidistributed points on the circle in the mean-field limit, and stays near it for long time. This is done by a study of the linearized system and leveraging nonlinear stability results from [grenier2000nonlinear]. See also [bruno2025multiscale, karagodin2024clustering, alcalde2025attention].
Our interest lies in the number of clusters during the first metastable phase in dimension $d=2$. At time $\tau=0$, we initialize $n$ tokens at iid uniform points on the circle. Under the self-attention flow, tokens follow the vector field $\nabla\log\big(K_{\beta^{-1/2}}*\mu_{\tau}\big)$, so metastable clusters coincide with local maxima of the smoothed empirical measure $K_{\beta^{-1/2}}*\mu_{\tau}$. In particular, at early times the circle is partitioned by the stationary points of $K_{\beta^{-1/2}}*\mu_{0}$, and each arc contracts toward its nearest maximum, making the number of clusters equal to the number of these maxima. Our $1d$ analysis shows that, for iid Gaussian data, the maxima concentrate in two symmetric belts where $t^{2}\in[2\log n-3\log\beta,2\log n-\log\beta]$; on the circle this corresponds to angular locations where the local inter-token spacing is $\simeq\beta^{-\frac{1}{2}}$, explaining both the $\sqrt{\beta}$ scaling of the metastable cluster count and their preferred positions.
Here, we focus on a simplified setting by working on the real line instead of the circle (or higher-dimensional spheres), but we believe the analysis could be extended to these cases pending technical adaptations. Notwithstanding, Theorem˜1.2 reflects what is seen in simulations (Figure˜3).
1.3 Sketch of the proof
The spirit of the proof of results such as Theorem˜1.1 and others presented in [mammen91, mammen95, mammen97] is similar to ours—one applies the Kac-Rice formula (Theorem˜2.1) to a Gaussian approximation of $(\widehat{P}_{n}^{\prime}(t),\widehat{P}_{n}^{{}^{\prime\prime}}(t))$ and argues its validity. However, the main limitation of these works is that modes are counted in a fixed and finite interval $[a,b]$ (and $0,1]^{d}$ in the higher dimensional cases). Extending these techniques to the whole real line demands for different, significantly stronger, approximation results using Edgeworth expansions.
We sketch the key ideas that allow us to count modes over the real line. We truncate $\mathbb{R}$ to the interval
@@ 1.3 Sketch of the proof
and so $T$ is well-defined for large $\beta$. Motivated by Theorems˜1.1 and 1.2, we also define the interval
@@
if $\beta\leqslant n^{\frac{2}{3}}$ and define $T^{\prime}=\varnothing$ if $\beta>n^{\frac{2}{3}}$. We use the Kac-Rice formula to compute the expected number of modes of $\widehat{P}_{n}$ in the symmetric intervals $T$ and $T^{\prime}$. All asymptotics are as $n,\beta\to\infty$.
@@
    In expectation over $X_{i}$, the number of modes of $\widehat{P}_{n}$ in $T^{\prime}$ is $O\mathopen{}\mathclose{{\left(\sqrt{\beta}}}\right)$.
The Kac-Rice computation appears tractable only when the joint distribution of $(\widehat{P}_{n}^{\prime}(t),\widehat{P}_{n}^{\prime\prime}(t))$ is Gaussian, which it is not. To overcome this obstacle, we apply the Kac-Rice formula over a Gaussian approximation of the joint distribution in Section˜2. For the specific underlying density and KDE in ˜1.1, we are able to justify in Section˜3 the approximation for all $t$ in the growing interval $T$ instead of a fixed interval. This is why Theorem˜1.1 only counts modes in a fixed interval.
To show the validity of the Gaussian approximation, we use the Edgeworth expansion of the joint distribution of $(\widehat{P}_{n}^{\prime}(t),\widehat{P}_{n}^{\prime\prime}(t))$ around the Gaussian distribution with matching first two moments. We bound the error due to the third order term of the expansion directly, and deal with the higher order terms by appealing to the error bounds of densities in the Edgeworth approximation similar to [bhattacharya2010normal, Theorem 19.2 and 19.3]. This strategy has been used in [bally2019non], but in a completely different context. We note that [mammen97] employ the same theorem to justify the Gaussian process approximation over a fixed interval.
Indeed, as $T,\beta$ grow with $n$, the error decay rate of the Edgeworth approximation is quite delicate near the boundary of $T$. Instead of the usual case of powers of $n^{-\frac{1}{2}}$, it is powers of $e^{-\frac{\omega(\beta)}{4}}$ (see Lemma˜3.5 and ˜3.6). This is exactly why we need to introduce the $\omega(\beta)$ term in $T$. In doing so, we will see that the normal approximation is invalid outside of $T$ (see Remark˜3.4), but crucially $T$ is sufficiently large to cover almost all modes, as observed empirically in Remark˜1.4 and Figure˜4 and given below.
@@
If $n^{c}\lesssim\beta\lesssim n^{2-c}$ for arbitrarily small $c>0$, then the expectation over $X_{i}$ of the number of modes of $\widehat{P}_{n}$ that lie outside of $T$ is $O\mathopen{}\mathclose{{\left(e^{\frac{\omega(\beta)}{2}}\sqrt{\beta}}}\right)$.
We prove this in Section˜4.2 with an argument from scale-space theory: we bound the number of modes outside $T$ by the number of samples $X_{i}$ outside $T$, which we then bound naively. This is precisely the argument used by [carreira2003number, Theorem 2] to show Gaussian mixtures over $\mathbb{R}$ with $n$ components must have at most $n$ modes. This argument crucially relies on the kernel density estimator being Gaussian (see Remark˜4.4).
Now, Theorem˜1.2 follows from Propositions˜1.7 and 1.8 provided $1\ll\omega(\beta)\ll\log\log\beta$. Indeed, the bounds on $\omega(\beta)$ are chosen to balance these error terms. In fact, all error terms other than the Kac-Rice integral over $T\setminus T^{\prime}$ of the Gaussian approximation of the density are $O\mathopen{}\mathclose{{\left(e^{-\frac{\omega(\beta)}{4}}\sqrt{\beta\log\beta}}}\right)$.